Suppose we are told that the acceleration (a) of a particle moving with uniform speed (v)


in a circle of radius (r) is a=kr^ n v^ prime


Determine the values of n and m and write the simplest form of this equation.


Note: K is a dimensionless constant


"Show your work"


1
Expert's answer
2021-12-02T10:05:47-0500

According to the principle of homogeneity in any physical formula, the dimensions of each term are the same.

Dimension of acceleration

"a = L\u2019T^{-2}"

Dimension of velocity

"V=LT^{-1}"

Given, that acceleration of the particle moving along path is

"a=kr^nv^m

=L^n(LT^{-1})^m = L^{m+m} T^{-m} \\

a =L’T^{-2} =L^{n+m} T^{-m}"

Then, according to the principle of homogeneity, this dimensional equation is balanced under the conditions

"1 = n+m \\

-2= -m \\

m=2 \\

n= 1-m = 1-2 = -1"

Then, we can write the acceleration expression as

"a=kr^nv^m = kr^{-1}v^2 =k\frac{v^2}{r}"

For uniform circular motion, k=1.

Then simplest form of the equation is,

"a = \frac{v^2}{r}"